L{ s/ (s2+1)2 }

1
Expert's answer
2021-09-05T18:29:08-0400
"f(t)=L^{-1}\big(\dfrac{s}{(s^2+1)^2}\big)=\dfrac{1}{2}L^{-1}\big(\dfrac{2(1)s}{(s^2+(1)^2)^2}\big)"

"=\dfrac{1}{2}t\sin t"


"f(t)=L^{-1}\big(\dfrac{s}{(s^2+1)^2}\big)=\dfrac{1}{2}t\sin t"